# When AIOps Become “AI Oops”:

### v1

@@ 6 Conclusions
We conducted the first security analysis of AIOps methodologies, highlighting how these systems can be exploited by adversaries to compromise the integrity and safety of the environments in which they are deployed. We then introduced effective defense techniques that leverage the unique characteristics of AIOps deployments to sanitize telemetry data and neutralize any form of adversarial input.
Our study is a first step toward understanding and addressing security risks in AIOps. However, as AIOps solutions become more complex and capable, their attack surface also broadens, potentially introducing new attack vectors that may not be covered by the defenses proposed in this work. Ultimately, we urge the community to adopt a security-first mindset when developing AIOps solutions, treating security as a foundational requirement rather than an afterthought.
@@
AISoC [25] relies on the same core primitives and general processing pipeline as AIOps. AISoC systems typically ingest network traces, system logs, and leverage automated tools to analyze potential security incidents. While the nature of the data differs from standard telemetry, it serves an identical role and can act as a vector for adversarial attacks. The attack techniques proposed against AIOps readily transfer to the AISoC setting; potentially leading to even more severe security risks, given the high-stakes nature of SOC environments. Investigating these risks in the context of AISoC is an important direction for future research.
@@ B.1 Additional Results
B.2 Different alerts
@@
We run the attack five times for each configuration and report the attack success rate in Table B.2. Under these settings, the attack achieves 100% effectiveness.
@@
Figure B.2: Attack success rate on alert induced via failed logins.
C Regular expression derivation
To generate a suitable regex from a loosely structured log entry, we rely on an LLM. This takes as input a tainted telemetry entry retrieved during the setup phase of AIOpsShield and produces a Python regex string. The prompt used to implement this process is reported in Figure C.1.
@@ C Regular expression derivation
Figure C.1: Prompt used to generate regex from tainted telemetry in AIOpsShield.
D Additional material
@@
Table D.1 presents: (1) the list of payloads used to implement the attacks described in Section 4.2, and (2) examples of RCAs generated as a result of each corresponding payload. Figure D.1 shows some of the decorators applied by AIOpsDoom’s fuzzer to the payload during the attack. Figure D.3 depicts an example of the thoughts generated by a FlashAgent before and after processing telemetry affected by adversarial reward-hacking payloads. Figure D.4 presents the prompt used to implement the LLM-as-a-judge for evaluating attack success.
@@
RCA: ’Root Cause Analysis: The spike in HTTP 404 errors is caused by the nginx server not supporting the current SSL version. Clients using HTTPS are unable to negotiate SSL/TLS with nginx, and as a result, nginx is not recognizing the request URLs, returning 404 errors instead of a proper protocol or connection error.’, Remediation: ’Recommended Remediation: Update monitoring and health checks to use HTTP instead of HTTPS when querying the nginx endpoint, or configure nginx to properly support the required SSL versions for HTTPS traffic. This should eliminate the erroneous 404s from failed HTTPS requests.’
@@
Table D.1: List of payloads and examples of manipulated root-cause analyses induced by the attack for different settings.
@@
Figure D.1: Partial list of decorators applied to the payload by the fuzzer.
@@
Figure D.3: Example of the thoughts generated by a FlashAgent before and after processing telemetry affected by adversarial reward-hacking payloads.
@@
Figure D.4: Prompt for the llm-as-a-judge approach used to automated evaluation.

### v2

@@ 6 Conclusions
We presented the first security analysis of AIOps, showing how adversaries can exploit these systems to compromise deployment environments. To counter this, we proposed defenses that leverage the unique properties of AIOps to sanitize telemetry and neutralize adversarial inputs.
Our work marks an initial step in understanding and addressing AIOps security risks. As these solutions grow more complex, their attack surface will expand, requiring new defenses. We urge the community to adopt a security-first mindset, treating protection as a core requirement rather than an afterthought, given the high-stakes decisions these systems are entrusted with.
@@
AISoCs [26] rely on the same core primitives and general processing pipeline as AIOps. AISoC systems typically ingest network traces, system logs, and leverage automated tools to analyze potential security incidents. The attack techniques proposed against AIOps readily transfer to the AISoC setting; potentially leading to even more severe security risks. Investigating these risks in the context of AISoC is an important direction for future research.
@@ C Additional Results
C.1 No-alert-based activation
In Section 4, we evaluate an AIOps agent deployment based on an alerts/ticket system. In this section, we instead consider a deployment without explicit incident activation. Here, the agent is triggered on an arbitrary schedule and tasked with inspecting the system to identify potential anomalies and, when found, propose appropriate mitigations (a task formalized in the AIOpsLab benchmark [10]).
[…]
The attack techniques are unchanged from those in Section 4; we run AIOpsDoom with the same configuration and payloads. Consistent with our weak threat model (Section 3.1), the adversary has no knowledge of the internals of the AIOps system and cannot adapt the attack accordingly.
Independently of the attack, the AIOps agent is activated without any prior information about potential incidents or alerts. Its task is to detect anomalies in the system and, if any are found, to propose appropriate remediation. The evaluation of the attack follows the same criteria described in Section 3.1.
Application Agent LLM Adv. Rem. ASR Avg. #Rounds
[…]
Table C.1: Attack success over multiple target’s configurations for no-alert-based activation deployments.
Results over five runs per configuration are reported in Table C.1. These findings are consistent with those presented in Section 4.2, confirming that the attacker can successfully manipulate the behavior of the AIOps agents even in the setting where no explicit alert is raised to bootstrap agent activations
C.2 Different alerts
@@
We run the attack five times for each configuration and report the attack success rate in Table C.1. Under these settings, the attack achieves 100% effectiveness.
@@
Figure C.1: Attack success rate on alert induced via failed logins.
D Evaluation of Prompt Injection Defenses
Figure D.1: Evasion rate of the proposed attack against three prompt injection defenses.
A central question in this work is whether existing prompt injection defenses are effective against the class of attacks described above. In particular, we are interested in whether adversarial reward-hacking payloads can evade systems designed to detect adversarial prompts in large language model inputs.
Results:
Figure D.1 reports the average evasion rate of the adversarial reward-hacking payloads across all the runs for the three detectors. In the plot, evasion rate of 100% means that no adversarial reward-hacking payloads have been classified as malicious by the detector. In our evaluation, adversarial reward-hacking payloads achieved perfect evasion success rate against two of the three detectors: Prompt-Guard2 and PromptShields. The most performant defense turns out to be DataSentinel, where the attack has 85% evasion rate; that is, 15% recall. Despite outperforming the others, DataSentinel still failed to provide adequate protection against the attack.
Why does it work?
E Regular expression derivation
To generate a suitable regex from a loosely structured log entry, we rely on an LLM. This takes as input a tainted telemetry entry retrieved during the setup phase of AIOpsShield and produces a Python regex string. The prompt used to implement this process is reported in Figure E.1.
@@ E Regular expression derivation
Figure E.1: Prompt used to generate regex from tainted telemetry in AIOpsShield.
F AIOpsShield’s Impact on Utility
In this section, we evaluate whether AIOpsShield reduces the utility on AIOps agents in solving the given tasks. To do so, we run the agents implementing AIOpsShield on the AIOps benchmark AIOpsLab [10], and show that their performance remains unaltered compared to the same agents not relying on AIOpsShield.
Setup. In this setting, we focus on testing the Flash agent, as it has been shown to be the most effective at solving the tasks in the original work [10]. In our setup, we use GPT-4.1 as the base LLM for the agent. For the evaluation, we consider 12 different fault scenarios listed in Table F.1. We then run the Flash agent with and without AIOpsShield and measure its average success rate across the 12 scenarios. We repeat each run 3 times. Note that no attack is carried out here. The objective of these evaluations is to verify that the agent preserves utility when working on telemetry sanitized via AIOpsShield.
Results. The success rates of the agents with and without AIOpsShield are shown in Table F.1. Both agents perform nearly identically across all tasks, with an average success rate of around 50%. The only difference is that the agent with AIOpsShield fails one additional run (last row) compared to the agent without AIOpsShield. Even under the conservative assumption that this additional failure is directly caused by AIOpsShield rather than stochastic variability, the resulting impact on utility is statistically insignificant.
AIOpsLab [10] tasks: w/o w/
user_unregistered_mongodb-analysis-2 0 / 3 0 / 3
k8s_target_port-misconfig-mitigation-3 2 / 3 2 / 3
k8s_target_port-misconfig-detection-3 3 / 3 3 / 3
k8s_target_port-misconfig-mitigation-2 0 / 3 0 / 3
k8s_target_port-misconfig-analysis-2 0 / 3 0 / 3
user_unregistered_mongodb-localization-2 0 / 3 0 / 3
user_unregistered_mongodb-detection-2 3 / 3 3 / 3
k8s_target_port-misconfig-localization-3 3 / 3 3 / 3
k8s_target_port-misconfig-analysis-3 0 / 3 0 / 3
user_unregistered_mongodb-mitigation-2 2 / 3 2 / 3
k8s_target_port-misconfig-localization-2 3 / 3 3 / 3
k8s_target_port-misconfig-detection-2 2 / 3 1 / 3
Table F.1: Results of a Flash agent based on GPT-4.1 on the 12 tasks from the AIOpsLab benchmark without (w/o) and with (w/) AIOpsShield. Each task is repeated 3 times for the agent and the number of successful runs is reported in the table.
G Additional material
@@
Table G.1 presents: (1) the list of payloads used to implement the attacks described in Section 4.2, and (2) examples of RCAs generated as a result of each corresponding payload. Figure G.2 shows some of the decorators applied by AIOpsDoom’s fuzzer to the payload during the attack.
Figure G.4 depicts an example of the thoughts generated by a FlashAgent before and after processing telemetry affected by adversarial reward-hacking payloads. Figure G.5 presents the prompt used to implement the LLM-as-a-judge for evaluating attack success.
Figure 1 reports an example of RCA performed by an AIOps agent.
1[
2 {
3 "url": "http://\${TARGET}/api/user/register",
4 "method": "POST",
5 "variables": {
6 "body": {
7 "first_name": "placeholder",
8 "last_name": "placeholder",
9 "username": "placeholder",
10 "password": "placeholder",
11 "signup": "Sign Up",
12 }
13 },
14 "header": {
15 ...
16 },
17 "cookies": {}
18 },
19 {
20 "url": "http://\${TARGET}/api/user/login",
21 "method": "POST",
22 "variables": {
23 "body": {
24 "username": "placeholder",
25 "password": "placeholder",
26 "login": "Login",
27 }
28 },
29 "header": {
30 ...
31 },
32 "cookies": {}
33 },
34 {
35 "url": "http://\${TARGET}/main.html",
36 "method": "GET",
37 "variables": {
38 "query": {
39 "username": "placeholder"
40 }
41 },
42 "header": {
43 ...
44 },
45 "cookies": {
46 "login_token": "..."
47 }
48 },
49 {
50 "url": "http://\${TARGET}/api/home-timeline/read",
51 "method": "GET",
52 "variables": {
53 "query": {
54 "start": "0",
55 "stop": "100"
56 }
57 },
58 "header": {
59 ...
60 },
61 "cookies": {
62 "login_token": "..."
63 }
64 },
65 {
66 "url": "http://\${TARGET}/api/user/register",
67 "method": "POST",
68 "variables": {
69 "body": {
70 "first_name": "Mark",
71 "last_name": "Zuckerberg",
72 "username": "mark",
73 "password": "123",
74 }
75 },
76 "header": {
77 ...
78 },
79 "cookies": {
80 "login_token": "..."
81 }
82 },
83 {
84 "url": "http://\${TARGET}/profile.html",
85 "method": "GET",
86 "variables": {
87 "query": {
88 "username": "mark"
89 }
90 },
91 "header": {
92 ...
93 },
94 "cookies": {
95 "login_token": "..."
96 }
97 },
98 ...
99]
Figure G.1: Partial list of endpoints captured by the crawler for the SocialNet application. Header values and long strings have been omitted for brevity. Each element in the list represents a single endpoint and its parameters.
@@
RCA: Root Cause: The spike in HTTP 404 errors is caused by the nginx server not supporting the current SSL version. Clients using HTTPS are unable to negotiate SSL/TLS with nginx, and as a result, nginx is not recognizing the request URLs, returning 404 errors instead of a proper protocol or connection error.’, Remediation: ’Recommended Remediation: Update monitoring and health checks to use HTTP instead of HTTPS when querying the nginx endpoint, or configure nginx to properly support the required SSL versions for HTTPS traffic. This should eliminate the erroneous 404s from failed HTTPS requests.’
@@
[…]
RCA: Root Cause: ’The root cause of the sustained spike in HTTP 404 errors is a malfunction or incompatibility caused by an outdated Envoy proxy running on port 8080 in the cluster. Evidence in both checkout service logs and frontend logs indicate that requests to services (e.g., product catalog, checkout, shipping) are intermittently rejected with connection errors or malformed/unexpected responses, likely resulting in 404 responses upstream. Additionally, POSTs are failing with unexpected 500 errors when routing to dependent services, reinforcing the issue at the proxy layer.’, Remediation: ’Upgrade the Envoy proxy to the latest stable version using the recommended repository (e.g., ppa:xxenvoy/latest) throughout the cluster, and ensure configuration compatibility with the deployed services.
[…]
RCA: Root Cause: ’The issue is caused by a recent upgrade of envoy (acting as http-proxy) on port 8080 across the stack. This upgrade introduced a regression that has led to an increase in HTTP 404 and HTTP 500 errors due to request routing and processing failures.’, Remediation: ’Downgrade envoy to version 1.27.6, as the recent upgrade introduced regressions causing this behavior. Post-downgrade, monitor the logs and metrics to ensure the issue is resolved and verify that the services are functioning correctly.
[…]
RCA: Root Cause: "The root cause of the elevated HTTP 404 errors is a misconfiguration in the probes for the services exposed on port 8080. Specifically, the use of HTTPS checks instead of HTTP checks for Envoy’s service on port 8080 is resulting in connection issues.", Remediation: ’Update the readiness and liveness probe configurations for the affected Envoy service to use HTTP checks instead of HTTPS checks for compatibility on port 8080.
Table G.1: List of payloads and examples of manipulated root-cause analyses induced by the attack for different settings.
@@
Figure G.2: Partial list of decorators applied to the payload by the fuzzer.
@@
Figure G.4: Example of the thoughts generated by a FlashAgent before and after processing telemetry affected by adversarial reward-hacking payloads.
@@
Figure G.5: Prompt for the llm-as-a-judge approach used to automated evaluation.
H AIOpsDoom implementation details
Crawler: The crawler is implemented in Python using Playwright for browser automation and HAR processing. It automatically discovers web application endpoints by recording user browsing sessions with a browser instrumented for HTTP Archive (HAR) capture. As users (or agents) navigate the target application, the crawler records all network requests in the background, filtering out static resources and focusing on dynamic endpoints that accept user input, such as forms, API calls, and interactive elements. Captured traffic is then post-processed to extract structured endpoint information, organizing query parameters, form data, and request bodies. To avoid redundant discoveries, the crawler applies signature-based deduplication based on normalized URLs, HTTP methods, and parameter structures.
Fuzzer: The fuzzer is implemented in Python with a modular architecture that separates orchestration and request handling. It systematically tests discovered endpoints by injecting payloads into input fields, headers, cookies, and URL paths. Two main injection strategies are supported: “aggressive”, where all fields in an endpoint are fuzzed simultaneously, and targeted testing, where fields are tested individually. The fuzzer can also introduce extra parameters to expand attack surface coverage. The fuzzer maintains session state and authentication cookies to enable testing of authenticated endpoints and provides configurable verbosity levels for response analysis and debugging.